OpenStudy (anonymous):

you need 5 men that are monochromatic colorblind to participate in a study. what is the probability you have to sample 20 men to find these 5 that are color blind?

OpenStudy (anonymous):

\[\left(\begin{matrix}20 \\ 5\end{matrix}\right) .01^5 (1-.01)^15\]

OpenStudy (anonymous):

i got .000001 but my professor took off two points

OpenStudy (anonymous):

on the 20 choose 5 and i dont understand why

Miracrown (miracrown):

what is the chance that someone is colorblind?

OpenStudy (anonymous):

The question you answered was "If I choose 5 men from a group of 20, what's the probability that they are all color blind."

OpenStudy (anonymous):

1% @Miracrown

OpenStudy (anonymous):

ok so how could i have written it to say that 5 are blind from the 20?

OpenStudy (anonymous):

reverse?

OpenStudy (anonymous):

5 choose 20?

Miracrown (miracrown):

i don't think it's that straightforward

Miracrown (miracrown):

i'm thinking kinda like a geometric distribution

OpenStudy (anonymous):

if its the geometric then it would be \[\theta(1-\theta) ^{x-1}\]

OpenStudy (anonymous):

so it would be \[.01(1-.01)^{15}\]

OpenStudy (anonymous):

Well, that's the probability of finding 15 normal sighted before your first color-blind.

OpenStudy (anonymous):

so i have to change the 15?

OpenStudy (anonymous):

Oh, right. Ok. So, I don't remember the reason for this, but your binomial coefficient should be: \[\left(\begin{matrix}N-1 \\ k-1\end{matrix}\right)\]

OpenStudy (anonymous):

The reason has something to do with having k-1 successes before the Nth trial...or something.

OpenStudy (anonymous):

yes i recognize that equation

OpenStudy (anonymous):

Ah, right. The probability of getting 4 color blind in 19 tries, then the probability of the last one being color blind.

OpenStudy (anonymous):

so 19 choose 4?

OpenStudy (anonymous):

\[\left(\begin{matrix}N-1 \\ k-1\end{matrix}\right) p^k (1-p)^{N-k}\]

OpenStudy (anonymous):

ok let me solve it

OpenStudy (anonymous):

33.3359?

OpenStudy (anonymous):

I get something similar, but MUCH smaller.

OpenStudy (anonymous):

3.33359?

OpenStudy (anonymous):

I ended up with 3.33359 x 10^(-7)

OpenStudy (anonymous):

yes I got the same

OpenStudy (anonymous):

thank you so much! :)

OpenStudy (anonymous):

K. It was just that 10^-7 I was talking about. Same result, but smaller :P Good work!